OpenStudy (anonymous):

Can you explain in words&numbers on how to solve what is x? As stated: -31-4x=-5-5(1+5x) Please help! Thank you

OpenStudy (anonymous):

ok, so we start with -31-4x=-5-5(1+5x) // first multiply the 5 through on the right hand side -31-4x=-5-5-25x // then add the two -5's together on the right hand side of the equation. -31-4x=-10-25x //then add 25x to both sides removing the 25x from the right hand side and (25x-4x) = 21x on the left side. -31+21x=-10 //then add 31 to both sides of the equation removing the -31 from the left hand side. (-10+31) = 21 21x=21 //then divide both sides by 21 so that you have just x on the left hand side of the equation. x=1

OpenStudy (anonymous):

Thank you, that helped a lot! @tomo

OpenStudy (anonymous):

yw